### v2

# On the Girth of Graph Lifts

Abstract
The size of the smallest $k$-regular graph of girth at least $g$ is denoted by the well-studied function $n(k,g)$. We introduce an analogous function $n(H,g)$, defined as the smallest size graph of girth at least $g$ that is a lift (or cover) of the, possibly non-regular, graph $H$. We prove that the two main combinatorial bounds on n(k,g)n(k,g)— the Moore lower bound and the Erdös-Sachs upper bound—carry over to the new lift setting.
We also consider two other functions: i) The smallest size graph of girth at least $g$ sharing a universal cover with $H$. We prove that it is the same as $n(H,g)$ up to a multiplicative constant. ii) The smallest size graph of girth least $g$ with a prescribed degree distribution. We discuss this known generalization and argue that the new suggested definitions are superior.
We conclude with experimental results for a specific base graph, followed by conjectures and open problems for future research.
Keywords: Lifts of graphs, Graph cover, Girth, Moore bound, Universal cover, Perron eigenvalue, Non-regular graph, Growth rate, Greedy Algorithms.
1 Introduction
The girth of a graph is the length of its shortest cycle. It is a classical challenge to determine $n(k,g)$, the smallest possible number of vertices $n$ of a $k$-regular graph of girth at least $g$. Minimal size $k$-regular graphs of girth $g$ are called $(k,g)$-cages, see [3, 6] for a survey of the subject. There are two simple combinatorial arguments yielding lower and upper bounds on $n(k,g)$, namely the Moore lower bound and the Erdös-Sachs 1963 upper bound, [5]. Their asymptotic version is:
$\Omega({(k-1)}^{g/2})\leq n(k,g)\leq O({(k-1)}^{g})$ (1)
To this day, the Moore bound remains the best asymptotic lower bound, and Erdös-Sachs achieves the best asymptotic upper bound by means of a combinatorial construction. There are number theoretic constructions that do better, such as the celebrated 1988 construction by Lubotzky, Phillips and Sarnak [13], that improves the upper bound to $O({(k-1)}^{3g/4})$.
Less is known about the minimal size of non-regular girth $g$ graphs. The result of Alon, Hoory and Linial 2002 [1] generalizes the Moore bound to non-regular graphs. They proved that the smallest possible number of vertices of a girth $g$ graph with minimal degree at least two and degree distribution $\cal D$} is lower bounded by:
$n({\cal D},g)\geq\Omega(\Lambda({\cal D})^{g/2})$ (2)
where $\Lambda=\Lambda({\cal D})=\prod_{d}{(d-1)}^{(d\,p_{d})/{(\sum_{d^{\prime}}d^{\prime}p_{d^{\prime}}})}$ and $p_{d}$ is the portion of degree $d$ vertices specified by $\cal D$}. The base of the exponent, $\Lambda$, is some weighted average of the degree minus one, known to be lower bounded by the average degree minus one.
However, accepting $n({\cal D},g)$ as the right analogue of $n(d,g)$ to non-regular graphs has several drawbacks. Firstly, the Erdös-Sachs construction does not carry over to the new setting, so there is no ”trivial” upper bound. Secondly, the existence of Moore graphs, achieving tight equality with the non-asymptotic version of (2), does not seem possible, expect for degenerate cases, as indicated by Eisner and Hoory [4].
In this work, we argue that there is a better way to formulate the problem using lifts of graphs. Given any finite connected graph $H$ with minimal degree at least two, $n(H,g)$ is the smallest lift of $H$ with girth at least $g$. We prove that the Moore bound and the Erdös-Sachs construction naturally generalize to the new setting, even in their non-asymptotic form. For the sake of avoiding complications, we exclude the case of $H$ being a cycle, as it is well understood. We also note that the $n(H,g)$ problem is not a generalization of the $n(k,g)$ problem, since the family of all $k$-regular graphs is not necessarily a lift of some base graph $H$.
[…]
Theorem 1.1.
For a finite connected multigraph with loops $H$ with minimal degree at least two and maximal degree greater than two:
$\Omega({\rho(\tilde{H})}^{g/2})\leq n(H,g)\leq O({\rho(\tilde{H})}^{g}),$ (3)
[…]
2 Lifts of Graphs and the Growth Rate of Balls
Graphs are always undirected and connected, with minimal degree $\geq 2$ and maximal degree >2>2. The graphs are not necessarily simple, as we allow multiple edges and two types of self loops: half-loops, and whole-loops, following the convention of Friedman and others [2, 7, 8, 11]. If the graph is simple, the extra notation can be ignored, and each edge is regarded as a pair of directed edges, each being the inverse of the other.
A graph $H$ is the tuple $(V,\overline{E},h,t,\iota)$, where $V$ is the vertex set, $\overline{E}$ is the directed edge set, $h,t:\overline{E}\rightarrow V$ respectively map an edge to its head or tail vertex, and the inverse $\iota$, maps an edge to its inverse $\iota(e)=e^{-1}$. The obvious relations hold: $h(e^{-1})=t(e)$ and that the inverse mapping is an involution. Note that while a half-loop contributes one to the degree, a whole-loop contributes two. An edge $e$ with $h(e)=t(e)$ is called a loop: a half-loop if $e=e^{-1}$ and a whole-loop otherwise. The degree of $v$, denoted $\deg(v)$ is the number of edges whose tail is $v$, i.e. $\deg(v)=|t^{-1}(v)|=|h^{-1}(v)|$. The neighborhood of $v$ is $\{h(e):e\in t^{-1}(v)\}$, namely the heads of edges emanating from $v$. A length $k$ walk $\omega$ from $u$ to $v$ is an edge sequence $e_{1},\ldots,e_{k}$ such that $t(e_{1})=u$, $h(e_{k})=v$ and $h(e_{i})=t(e_{i+1})$ for all $i$. The walk $\omega$ is non-backtracking if $e_{i+1}\neq(e_{i})^{-1}$ for all $i$. Let $\mbox{dist}(u,v)$ be the length of the shortest walk from $u$ to $v$ (necessarily, such a walk is always non-backtracking). Denote the radius $r$ ball centered at some vertex $v$ or edge $e$ by $B_{r}(v)=\{u:\mbox{dist}(u,v)\leq r\}$ and $B_{r}(e)=B_{r}(h(e))\cup B_{r}(t(e))$.
Let $H$ be a finite undirected and connected graph, as above. Following the definition of a topological covering we say that the graph $G$ covers $H$, or that $G$ is a lift of $H$, if there is a surjective homomorphism $\pi:G\rightarrow H$ that is a bijection of the edges emanating from $v$ to the edges emanating from $\pi(v)$ for all $v\in V(G)$, i.e. $\pi(t^{-1}(v))=t^{-1}(\pi(v))$. One can verify that fiber size is fixed for all vertices and edges. Namely, there is an integer $n$, which we call the height of the lift, such that $|\pi^{-1}(v)|=|\pi^{-1}(e)|=n$ for all $v\in V(G)$ and $e\in\overline{E}$. We identify $V(G)$ with $V(H)\times[n]$ and $\overline{E}(G)$ with $\overline{E}(H)\times[n]$, where $n]=\{1,2,\ldots,n\}$. Furthermore, we regard $\pi$ as a mapping $\mbox{perm}_{\pi}$ from $\overline{E}(H)$ to the permutation group $S_{n}$, where for any $e\in\overline{E}(H)$ and $i\in[n]$ the edge in $\pi^{-1}(e)$ with tail $(t(e),i)$ has its head at $(h(e),\mbox{perm}_{\pi}(e)(i))$. It follows that $\mbox{perm}_{\pi}(e^{-1})=\mbox{perm}_{\pi}(e)^{-1}$ and that $\mbox{perm}_{\pi}$ of a half-loop is an involution.
The universal cover of $H$, denoted by $\tilde{H}$, is the graph covering all lifts of $H$. This graph, which is unique up to isomorphism, is an infinite tree. One way to construct $\tilde{H}$ is to choose some arbitrary vertex $v_{0}\in V(H)$ and let V(H~)V($\tilde${H}) be the set of all finite non-backtracking walks from $v_{0}$. Two vertices of $\tilde{H}$ are adjacent if one walk extends the other by a single step and the root of the tree is the empty walk $\epsilon$.
We define the growth rate of $\tilde{H}$ by $\rho(\tilde{H})=\lim_{r}|B_{r}(v)|^{1/r}$, where $v$ is some arbitrary vertex in V(H~)V($\tilde${H}). The non-backtracking adjacency operator of $H$ is the 0-1 matrix $B$ of directed edges versus directed edges, where $B_{f,e}=1$ if $t(f)=h(e)$ and $e\neq f^{-1}$. Then by the following propositions, as $H$ has a minimal degree of $\geq 2$ and maximal degree >2>2, the matrix $B$ is non-negative irreducible, so by the Perron-Frobenius Theorem, its spectral radius $\rho(B)$ is an algebraically simple eigenvalue corresponding to strictly positive left and right eigenvectors. Furthermore, all eigenvalues whose absolute value is equal to $\rho(B)$ are algebraically simple.
Proposition 2.1 ([10] proposition 3.3,[4]).
The non-backtracking adjacency matrix $B_{H}$ is irreducible iff $H$ is connected with $\operatorname{mindeg}(H)\geq 2$ and $\operatorname{maxdeg}(H)>2$.
Proposition 2.2 (Perron–Frobenius, Theorem 8.4.4 from [9]).
Given a non-negative irreducible matrix $M\in M_{n}(\mathbb{R})$ for $n\geq 2$, the spectral radius $\rho(M)$ is a positive and algebraically simple eigenvalue. Furthermore, up to scaling there are unique vectors $\mathbf{x},\mathbf{y}$ so that $M\mathbf{x}=\rho(M)\mathbf{x}$ and $\mathbf{y}^{T}M=\rho(M)\mathbf{y}^{T}$; these vectors are strictly positive.
Proposition 2.3 (Corollary 8.4.6 (c) from [9]).
Given a non-negative irreducible matrix $M\in M_{n}(\mathbb{R})$, let $k$ be the number of distinct maximal eigenvalues in absolute value. Then, these eigenvalues are algebraically simple and their values are $\rho(M)e^{2\pi ij/k}$ for $j=0,\ldots,k-1$.
Theorem 2.4.
[…]
$c_{1}\rho(B_{H})^{r}\leq|B_{r}(v)|\leq c_{2}\rho(B_{H})^{r}.$
Corollary 2.5.
For any finite undirected connected graph $H$ with $\operatorname{mindeg}(H)\geq 2$ and $\operatorname{maxdeg}(H)>2$, the growth rate of its universal cover $\tilde{H}$ is equal to the spectral radius of $B_{H}$:
$\rho(\tilde{H})=\rho(B_{H}).$
We provide two distinct proofs for Theorem 2.4. The first utilizes the algebraic structure of the non-backtracking matrix $B_{H}$, relying on its spectral properties as described in Proposition 2.3. The second employs combinatorial counting methods.
[…]
Proof I of Theorem 2.4.
Let $H$ be a graph satisfying the hypotheses, let $\tilde{H}$ be its universal cover, and let $\pi$ be the respective cover map. Let $v$ be an arbitrary vertex in $\tilde{H}$. Then, the size of the radius $r$ ball centered at $v$, is
$|B_{r}(v)|=1+\sum_{i=1}^{r}\Delta_{i},$ (4)
where $\Delta_{i}=\Delta_{i}(v)$ denotes the number of length $i$ non-backtracking walks from $v$. The quantity $\Delta_{i}$ can be expressed using the non-backtracking matrix $B_{H}$ as
$\Delta_{i}=\delta_{v}^{T}(B_{H})^{i-1}\mathbf{1},$ (5)
where $\mathbf{1}$ is the all-one vector indexed by the set of directed edges $\overline{E}$, and $\delta_{v}$ is the indicator vector on the edges emanating from $\pi(v)$, where the entry $(\delta_{v})_{e}$ is one if $t(e)=\pi(v)$ and zero otherwise.
Let $B_{H}=V^{-1}\left[\bigoplus_{j=1}^{m}J_{j}\right]V$ be the Jordan decomposition of $B_{H}$. Since $B_{H}$ is non-negative and irreducible, by Proposition 2.1, it follows by Proposition 2.3 that the maximum-modulus eigenvalues correspond to $k$ scalar $1\times 1$ Jordan blocks. Thus, the decomposition can be rearranged so that the Jordan blocks $J_{p}$ for $p=1,\ldots,k$ are the scalars $J_{p}=(\rho\,\omega_{k}^{p-1})$, where $\rho=\rho(B_{H})$ and $\omega_{k}=e^{2\pi i/k}$ (here $i$ denotes $\sqrt{-1}$).
Moreover, the left and right Perron eigenvectors of $B_{H}$, satisfying $\mathbf{y}^{T}B_{H}=\rho\mathbf{y}^{T}$ and $B_{H}\mathbf{x}=\rho\mathbf{x}$, are the first row of $V$ and first column of $V^{-1}$, i.e. $\mathbf{y}^{T}=(V)_{1,\cdot}$ and $\mathbf{x}=(V^{-1})_{\cdot,1}$.
Denoting the zero padded $J_{j}$ blocks by $\widehat{J_{j}}$, we write $B_{H}$ as
$B_{H}=V^{-1}\left[\sum_{j=1}^{m}\widehat{J_{m}}\right]V=V^{-1}\left[\sum_{p=1}^{k}E_{p}\rho\,\omega_{k}^{p-1}+R\right]V,$
where the matrix $E_{p}$ is all-zero except for a one in its $p$-th diagonal entry, and the residue matrix $R=\sum_{j=p+1}^{m}\widehat{J_{j}}$ satisfies the condition $\rho^{\prime}=\rho(R)<\rho$.
Applying this decomposition to the expression for $\Delta_{i}$ yields:
$\Delta_{i}$ $=\rho^{i-1}\sum_{p=1}^{k}(\delta_{v}^{T}V^{-1}E_{p}V\mathbf{1})\,\omega_{k}^{(p-1)(i-1)}+\delta_{v}^{T}V^{-1}R^{i-1}V\mathbf{1}$
$=\rho^{i-1}\sum_{p=1}^{k}a_{p}\,\omega_{k}^{(p-1)(i-1)}+r_{i},$
where $a_{p}=\delta_{v}^{T}V^{-1}E_{p}V\mathbf{1}$ and the residue term $r_{i}=\delta_{v}^{T}V^{-1}R^{i-1}V\mathbf{1}$.
The residue term $r_{i}$ is bounded by:
$|r_{i}|\leq\|\delta_{v}\|\rho(V^{-1})\rho(R)^{i-1}\rho(V)\|\mathbf{1}\|=\Big(\|\delta_{v}\|\rho(V^{-1})\rho(V)\Big)(\rho^{\prime})^{i}=o(\rho^{i})\,\,\text{ as }i\rightarrow\infty.$
Furthermore, $a_{p}$ are complex constants, and the coefficient $a_{1}$, corresponding to the Perron eigenvalue $\rho$, is positive:
$a_{1}=\delta_{v}^{T}V^{-1}E_{1}V\mathbf{1}=\delta_{v}^{T}(V^{-1}E_{1})(E_{1}V)\mathbf{1}=\delta_{v}^{T}\mathbf{x}\mathbf{y}^{T}\mathbf{1}>0.$
Now, let $i-1\equiv s\pmod{p}$, for $s\in\{0,\ldots,k-1\}$. Then, we can rewrite $\Delta_{i}$ as :
$\Delta_{i}=\rho^{i-1}\sum_{p=1}^{k}a_{p}\,\omega_{k}^{(p-1)s}+r_{i}=\rho^{i-1}\widehat{a_{s}}+r_{i},$
where $\widehat{a_{s}}=\sum_{p=1}^{k}a_{p}\,\omega_{k}^{(p-1)s}$. The sequence $\{\widehat{a_{s}}\}_{s=0,\ldots,k-1}$ is not all zero, because the coefficients $\{a_{p}\}_{p=1\ldots k}$ are not all zero and the matrix $\{\omega_{k}^{(p-1)s}\}_{p,s}$ is non-singular, being a Vandermonde matrix (specifically, the discrete Fourier transform matrix).
Since $|r_{i}|=o(\rho^{i})$, and since there exists at least one value $s_{0}$ for which $\widehat{a_{s_{0}}}\neq 0$, it follows that $\Delta_{i}=\Theta(\rho^{i})$ for all $i$ with $i-1\equiv s_{0}\pmod{p}$. Furthermore, as $\Delta_{i}$ is a monotone non-decreasing function of $i$, it follows that $\Delta_{i}=\Theta(\rho^{i})$ for all $i$.
Finally, as the assumptions on $H$ imply that $\rho>1$, the sum in (4) is asymptotically dominated by its last term, $\Delta_{r}$, implying that $|B_{r}(v)|=\Theta(\rho(B_{H})^{r}$ as $r\rightarrow\infty$, as claimed. ∎
Proof II of Theorem 2.4.
As in the first proof, let $H$ be a graph satisfying the hypotheses, let $\tilde{H}$ be its universal cover, let $\pi$ be the respective cover map and let $v$ be an arbitrary vertex in $\tilde{H}$. The non-backtracking adjacency matrix $B_{H}$ is non-negative irreducible by Proposition 2.1, implying that $\rho=\rho(B_{H})$ is an eigenvalue of $B_{H}$ with strictly positive left eigenvector $\mathbf{y}$, satisfying $\mathbf{y}^{T}B_{H}=\rho\,\mathbf{y}^{T}$. By (4) and (5) we express the size of the radius $r$ ball centered at $v$ by
$|B_{r}(v)|=1+\sum_{i=1}^{r}\mathbf{\delta_{v}}^{T}(B_{H})^{i-1}\mathbf{1},$
where $\mathbf{1}$ is the all-one vector indexed by the set of directed edges $\overline{E}$, and $\mathbf{\delta_{v}}$ is the indicator vector on the edges emanating from $\pi(v)$, where the entry $(\mathbf{\delta_{v}})_{e}$ is one if $t(e)=\pi(v)$ and zero otherwise. Also, as $B_{H}$is non-negative irreducible, all edges in $\overline{E}$ are reachable from $v$, implying that for some integer $r_{0}$ the following vector is strictly positive:
$\mathbf{z}^{T}=\sum_{i=1}^{r_{0}}\mathbf{\delta_{v}}^{T}(B_{H})^{i-1}.$
It follows that there is a constant $c_{1}>0$ such that $c_{1}\mathbf{y}\leq\mathbf{z}$ and therefore for all $r\geq r_{0}$ we can lower-bound the size of the radius $r$ ball centered at $v$ by
$|B_{r}(v)|\geq\sum_{i=r-r_{0}+1}^{r}\mathbf{\delta_{v}}^{T}(B_{H})^{i-1}\mathbf{1}=\mathbf{z}^{T}(B_{H})^{r-r_{0}}\mathbf{1}\geq c_{1}\rho^{r-r_{0}}\mathbf{y}^{T}\mathbf{1}\geq\Omega(\rho^{r}).\\$
As for the upper bound, as $\mathbf{y}$ is positive, there is a constant $c_{2}>0$ so that $\mathbf{\delta_{v}}\leq c_{2}\mathbf{y}$, and therefore
$|B_{r}(v)|\leq 1+c_{2}\sum_{i=1}^{r}\mathbf{y}^{T}(B_{H})^{i-1}\mathbf{1}=1+c_{2}(\mathbf{y}^{T}\mathbf{1})\sum_{i=1}^{r}\rho^{i-1}.$
By the hypotheses on $H$ we have $\rho=\rho(B_{H})>1$, implying that the geometric sum is $O(\rho^{r})$. ∎
3 Moore and Erdös-Sachs Bounds for Graph Lifts
In this section we prove the non-asymptotic form of Theorem 1.1 by means of Lemma 3.1 for the lower bound and Corollary 3.3 for the upper bound. The asymptotic result of Theorem 1.1 follows by combining the non-asymptotic results with Theorem 2.4, stating that the size of the radius $r$ ball in the universal cover grows as $\Theta(\rho^{r})$.
Lemma 3.1 (Moore bound for lifts).
Let $H$ be a finite undirected connected graph of minimal degree at least two. Then:
  • $n(H,g)\geq\max_{v\in V(\tilde{H})}|B_{\lfloor g/2\rfloor}(v)|$ if $g$ is odd
  • $n(H,g)\geq\max_{e\in E(\tilde{H})}|B_{\lfloor g/2\rfloor}(e)|$ if $g$ is even
Proof.
[…]
Lemma 3.2 (Erdös-Sachs bound for lifts).
Let $H$ be a finite undirected connected graph of minimal degree at least two and let $T$ be a spanning tree of $H$. Then for any integer $g\geq 2\mbox{diam}(T)+2$ there exists a graph $G$ covering $H$ with $\mbox{girth}(G)\geq g$ and $\mbox{diam}(G)\leq g+2\mbox{diam}(T)$.
Corollary 3.3.
Let $H$ be a finite undirected connected graph of minimal degree at least two, let $T$ be some spanning tree of $H$. Then for any integer $g\geq 2\mbox{diam}(T)+2$:
$n(H,g)\leq\min_{v\in V(\tilde{H})}|B_{g\,+\,2\mbox{diam}(T)}(v)|.$
Proof of Lemma 3.2.
Given some base graph $H$ and spanning tree $T$ satisfying the requirements, denote $g_{0}=2\mbox{diam}(T)+2$ and $D_{0}=g+2\mbox{diam}(T)$. The proof consists of two parts. The first part is to prove the existence of some graph $G$, possibly very large, covering $H$ with girth at least $g$. The second part is showing that as long as $\mbox{diam}(G)>D_{0}$, one can be trim $G$ into a smaller graph covering $H$ without introducing a cycle shorter than $g$. Repeating this procedure until $\mbox{diam}(G)\leq D_{0}$ yields the required result. We prove the existence of an initial graph $G$ separately in Lemma 3.4, and concentrate here of the trimming procedure. Throughout, we assume that $G$ is connected, as otherwise it can be replaced by one of its connected components.
Let $G$ be be some graph of girth $g\geq g_{0}$ covering $H$ with the cover map $\pi:G\rightarrow H$ with vertices $u^{\prime},v^{\prime}\in V(G)$ such that $\mbox{dist}(u^{\prime},v^{\prime})>D_{0}$. Let $u=\pi(u^{\prime})$, $v=\pi(v^{\prime})$, where possibly $u,v$ are the same vertex. Without loss of generality we identify $V(G)$ with $V(H)\times[n]$ and $E(G)$ with $E(H)\times[n]$ so that the $n$ copies of $T$ in $\pi^{-1}(T)$ respect layers. Namely, that for any edge $e\in E(T)$, both endpoint of each edge in $\pi^{-1}(e)$ are in the same layer. Furthermore, since $\mbox{dist}(v^{\prime},u^{\prime})>D_{0}>\mbox{diam}(T)$ it is clear than $v^{\prime}$ and $u^{\prime}$ are not in the same layer, and we may assume that $v^{\prime}=(v,n)$ and $u^{\prime}=(u,n-1)$.
[…]
Figure 1: Deleting two layers and adding edges for deficient vertices: (a) $e=(x,y)$ is not a half-loop, (b) $e=(z,z)$ is a half loop.
Figure 1 depicts the red edges above some non-tree edge in $e\in E(H)\setminus E(T)$ and its inverse. In case (a) $e=(x,y)$ is not its own inverse, and therefore has 4 distinct red edges, two above $e$ and two above $e^{-1}$. In case (b) $e=(z,z)$ is equal to its inverse and it has only two red edges.
The graph $G^{\prime\prime}$ is created from $G^{\prime}$ by adding the blue edges to account for the deficient vertices, as showed in Figure 1. For every $e\in E(H)\setminus E(T)$ we add the following two blue edges: $((t(e),m(e^{-1})),(h(e),l(e)))$, $((t(e),l(e^{-1})),(h(e),m(e)))$. In case (a) the blue edges for $e^{-1}$ are the inverse of the blue edges for $e$, while in case (b) the blue edges for $e$ are already an edge and its inverse.
We claim that $G^{\prime\prime}$ both covers $H$ and has girth at least $g$. The first claim easily follows by defining the map $\pi_{G^{\prime\prime}}:G^{\prime\prime}\rightarrow H$ to be $\pi_{G^{\prime}}$ with the newly added blue edges and checking it is a cover map.
In order to prove that $\mbox{girth}(G^{\prime\prime})$ is at least $g$, assume to the contrary that $G^{\prime\prime}$ has a length $k<g$ cycle $C$ with the edges $e_{1},e_{2},\ldots,e_{k}$. Since $\mbox{girth}(G^{\prime})\geq g$, the cycle $C$ has at least one new (blue) edge. We say that the new edge $e$ is positive if $t(e)$ is connected in $G$ by a red edge to $L_{n-1}$ and $h(e)$ is connected to $L_{n}$. Otherwise, $t(e)$ is connected to $L_{n}$ and $h(e)$ to $L_{n-1}$ and we say that $e$ is negative.
Figure 2: The cycle $C$ with new edges depicted in blue, along with their sign, and old paths existing in $G$ depicted by a black dotted line. (a) All new edges have the same sign, $P$ is some old path connecting adjacent blue edges, (b) Not all new edges have the same sign, there are at least two old paths ($P_{1}$ and $P_{2}$ in this case) connecting different sign blue edges.
If all new edges in $C$ have the same sign, as seen in Figure 2 (a), there is a path $P$ in $C$ of length at most k1k-1 that exists in $G$ and connects two same sign blue edges. Therefore, $P$ with two red edges forms a path connecting $L_{n}$ with $L_{n-1}$. It follows that $d_{G}(v^{\prime},u^{\prime})\leq 2\mbox{diam}(T)+k-1+2\leq 2\mbox{diam}(T)+g=D_{0}$, contradicting our assumption that $\mbox{dist}_{G}(v^{\prime},u^{\prime})>D_{0}$.
Otherwise, not all new edges have the same sign, as seen in Figure 2 (b). Therefore, there are at least two paths $P_{1}$ and $P_{2}$ in $C$, present in the original graph $G$, connecting two different sign new edges. Choosing $P$ to be the shortest of the two paths, yields a path $P$ of length at most $\lfloor(k-2)/2\rfloor$, with endpoints $h(e)$ and $t(f)$ for new edges $e,f$. Furthermore, both $h(e)$ and $t(f)$ are connected by red edges in $G$ to the same layer, either $L_{n}$ or $L_{n-1}$. Therefore, $G$ has a short cycle consisting of $P$, two red edges and a traversal of the tree $T$. The length of this cycle is at most $\lfloor(k-2)/2\rfloor+2+\mbox{diam}(T)\leq k/2+1+\mbox{diam}(T)<g/2+1+\mbox{diam}(T)$, which is less equal than $g$ for $g\geq g_{0}$, yielding a contradiction. ∎
Lemma 3.4.
For any undirected connected graph $H$ with minimal degree at least two and any $g\geq 1$, there exists a covering graph $G$ with $\mbox{girth}(G)\geq g$.
Proof.
Assume first that $H$ does not have half-loops. The proof is by induction on $g$, where the base $g=1$ trivially holds for $G=H$. By induction, assume that $G$ is a graph covering $H$ with $\mbox{girth}(G)\geq g-1$, where the number of length g1g-1 cycles $\phi(G)$ is as small as possible. If $\phi(G)=0$, then $\mbox{girth}(G)\geq g$, and we are done. Otherwise, assume $\phi(G)>0$ and let $G_{2}$ be a random 2-lift of $G$. Namely, a random graph with vertex set $V(G)\times\{0,1\}$, where for each edge $e\in\overline{E}(G)$ we pick the permutation $\mbox{perm}_{\pi}(e)\in S_{2}$ independently at random such that $\mbox{perm}_{\pi}(e)=\mbox{perm}_{\pi}(e^{-1})$. We observe that all 2-lifts of $G$ have girth at least g1g-1, and argue that among them there is a 2-lift with a smaller number of length g1g-1 cycles. Indeed, as every length g1g-1 cycle $C$ in $G$ is turned into either one cycle of length $2(g-1)\geq g$ or two cycles of length g1g-1, with equal probabilities, depending on the product of $\mbox{perm}_{\pi}(e)$ for $e\in C$, it follows that the expected value of $\phi(G_{2})$ is $\phi(G)$. However, since the lift with $\mbox{perm}_{\pi}(e)=\text{id}$ for all $e$ has $2\phi(G)$ length g1g-1 cycles, it follows that there exists some 2-lift $G^{\prime}$ of $G$ with $\phi(G^{\prime})<\phi(G)$, yielding a contradiction.
If $H$ has half-loops, we let the graph $H^{\prime}$ be some 2-lift of $H$, where the lift of the half-loop $e$ with $h(e)=t(e)=v$ is a single edge connecting the two vertices covering $v$. As the graph $H^{\prime}$ covers $H$ and does not have any half-loops, the required result is obtained by applying the preceding proof to $H^{\prime}$. ∎
4 Three Generalizations for Non-regular Graphs
Returning to our motivating problem of finding the right generalization of $n(d,g)$ to non-regular graphs, we propose three possible definitions. Let $n({\cal G},g)$ denote the size of the smallest girth $g$ graph in the graph family $\cal G$} and consider the following graph families for some fixed graph $H$:
${\cal L}(H)$ $=$ $\{G:G\mbox{ is a lift of }H\},$
${\cal U}(H)$ $=$ $\{G:\tilde{G}=\tilde{H}\}$,
${\cal D}(H)$ $=$ $\{G:G\mbox{ has the same degree distribution as }H\}.$
Respectively, $n(H,g)=n({\cal L}(H),g)$ is the smallest girth $g$ lift of $H$, $n({\cal U}(H),g)$ is the smallest girth $g$ graph with the same universal cover as $H$ and $n({\cal D}(H),g)$ is the smallest girth $g$ graph with the same degree distribution as $H$. Since ${\cal L}(H)\subseteq{\cal U}(H)\subseteq{\cal D}(H)$ it follows that:
$n({\cal L}(H),g)\geq n({\cal U}(H),g)\geq n({\cal D}(H),g).$ (6)
It turns out that the first inequality cannot be too far from equality:
Lemma 4.1.
For every finite, undirected and connected graph $H$, there exists a constant $c(H)$ such that:
$c(H)\cdot n({\cal U}(H),g)\geq n({\cal L}(H),g)\geq n({\cal U}(H),g).$
Proof.
As the right inequality is part of (6) we only need to prove the left inequality. Assume that $G$ is a girth $g$ graph with the same universal cover as $H$. By a result of Leighton [11], we know that for every two graphs $G_{1},G_{2}$ sharing the same universal cover, there is a graph covering both whose size is at most the product of their sizes times a constant depending only on the universal cover. Applying that result with $G_{1}=H$ and $G_{2}=G$ yields a graph $G^{\prime}$ covering both, as claimed. ∎
For the degree distribution based lower bound, we have the following result:
Proposition 4.2 (Alon, Hoory, Linial, [1]).
Let $G=(V,E)$ be an undirected girth $g$ graph on $n$ vertices with minimal degree at least two, and let:
$\Lambda=\prod_{v\in V}{(d_{v}-1)}^{d_{v}/|\overline{E}|},\;\;\;\;\;n_{0}(x+1,g)=\sum_{i=0}^{\lfloor(g-1)/2\rfloor}x^{i}+\sum_{i=0}^{\lfloor g/2-1\rfloor}x^{i}.$
Then:
  1. 1.
    $n\geq n_{0}(\Lambda+1,g)$:
  2. 2.
    $\Lambda\geq\overline{d}-1$, where $\overline{d}$ is the average degree of $G$, with equality only if $G$ is regular.
  3. 3.
    $\frac{1}{|\overline{E}|}\mathbf{1}^{T}(B_{H})^{r}\mathbf{1}\geq\Lambda^{r}$ for all $r\geq 0$.
It follows that:
$n({\cal D}(H),g)\geq n_{0}(\Lambda(H)+1,g)\geq n_{0}(\overline{d}(H),g).$ (7)
This result is related to the $\rho$ based lower bound of Theorem 1.1 by the following proposition. In that work, the authors prove that $\rho(H)\geq\Lambda(H)$ and specify two equivalent conditions for equality to hold.
Proposition 4.3 (Eisner, Hoory, [4]).
For a finite undirected and connected graph $H$ with minimal degree at least two and maximal degree greater than two, we have $\rho(H)\geq\Lambda(H)$. Furthermore, the following three conditions are equivalent:
  1. 1.
    $\rho(H)=\Lambda(H)$
  2. 2.
    For every non-backtracking cycle $C$ the geometric average of vertex degrees minus one along the cycle is equal to $\Lambda$,
    $\left(\prod_{v\in C}(\deg(v)-1)\right)^{\frac{1}{|C|}}=\Lambda(H).$
  3. 3.
    For every non-backtracking path $P$, where all its internal vertices (if any) have degree two and its endpoints have degree larger than two,
    $\left(\prod_{v\in P}(\deg(v)-1)\right)^{\frac{1}{2|P|}}=\Lambda(H).$
It follows from Claim 4.2, Proposition 4.3, inequalities (7), (6) and Theorem 1.1 that:
[…]
At this point, it is worthwhile to compare the values of $\rho$, $\Lambda$ and $\overline{d}$ for some specific graphs. Consider the three base graphs depicted in Figure 3:
Figure 3: (a) The bipartite $K_{3,2}$ graph, (b) $H_{23}$ with a degree 2 vertex $v$ and a degree 3 vertex $u$ with a half-loop edge, and (c) $K_{4}$ minus an edge, which is a 2-lift of $H_{23}$.
a. The complete bipartite graph $K_{3,2}$
  • $\overline{d}(K_{3,2})-1=(3+3+2+2+2)/5-1=1.4$
  • $\Lambda(K_{3,2})=2^{(3+3)/12}=\sqrt{2}$.
  • The universal cover of $K_{3,2}$ is the 3-regular infinite tree with a degree 2 vertex added to the middle of each edge. Therefore, $B_{r}(\cdot)$ has alternating layers of degree 2 and degree 3 vertices, as seen in Figure 4 (a), yielding $\rho(K_{3,2})=\sqrt{2}$. Note that $\rho$ can also be computed as the Perron eigenvalue of a 12 by 12 matrix, $B_{K_{3,2}}$ that can be reduced to a 2 by 2 matrix, as there are only two directed edge types:
    $\rho(K_{3,2})=\rho\left(\left[\begin{tabular}[]{cc}0&1\\ 2&0\end{tabular}\right]\right)=\sqrt{2}.$
b. The graph $H_{23}$
  • $\overline{d}(H_{23})-1=(3+2)/2-1=1.5$
  • $\Lambda(H_{23})=2^{3/5}=1.5157\ldots$.
  • As $H_{23}$ has 5 directed edges, $\rho(H_{23})$ is the Perron eigenvalue of the 5 by 5 matrix $B_{H_{23}}$. But, as there are only 3 directed edge types: $(u,v)$, $(u,u)$, $(v,u)$, it suffices to compute:
    $\rho(H_{23})=\rho(B_{H_{23}})=\rho\left(\left[\begin{tabular}[]{ccc}0&2&1\\ 0&0&1\\ 1&0&0\end{tabular}\right]\right)=1.5214\ldots$
c. The graph $K_{4}$ minus an edge
  • The graph is a 2-lift of $H_{23}$, so $\overline{d}$, $\Lambda$ and $\rho$ are the same.
It should be observed that because all three graphs are non-regular, we have strict inequality $\Lambda(H)>\overline{d}(H)-1$, in agreement with Claim 4.2. Furthermore, $\rho(H)>\Lambda(H)$, in agreement with Proposition [4]:
$\rho(K_{3,2})$ $=$ $\Lambda(K_{3,2})=\sqrt{2},$
$\rho(H_{23})=1.5214\ldots$ $>$ $\Lambda(H_{23})=2^{0.6}=1.5157\ldots.$
Figure 4: Two infinite trees: (a) $\widetilde{K_{3,2}}$ with $\rho=\sqrt{2}$, (b) $\widetilde{H_{23}}$ with $\rho=1.5214\ldots$.
The following diagram summarizes what we know about $H_{23}$:
[…]
5 Concrete Constructions and Bounds for the $H_{23}$ Graph
In this section we explore non-asymptotic upper and lower bounds on the value of $n(H,g)$, for the graph $H=H_{23}$ depicted in Figure 3 (b). We discuss this in the following subsections: The Moore lower bound, The Erdös-Sachs Upper Bound, Construction Algorithms, and Results.
5.1 The Moore Lower bound
The Moore lower bound for lifts is stated in Lemma 3.1, where the size of the ball centered around some tree vertex can be obtained by (4) and (5), with $g=2r+1$:
$n(H,g)\geq n_{0}(H,g)=\max_{v\in V(\tilde{H})}|B_{r}(v)|=1+\max_{v\in V(H)}\sum_{i=1}^{r}\delta_{v}^{T}\left(B_{H}\right)^{i-1}\mathbf{1},$ (10)
where $\delta_{v}$ is the indicator function of $t^{-1}(v)$. One can derive a similar formula for even girth.
5.2 The Erdös-Sachs Upper Bound
We compute the Erdös-Sachs upper bound for $n(H,g)$ using Corollary 3.3. As the spanning tree $T$ has diameter one, it follows that for all $g\geq g_{0}=4$:
$n(H,g)\leq\min_{v\in V(\tilde{H})}|B_{g+2}(v)|=1+\min_{v\in V(H)}\sum_{i=1}^{g+2}\delta_{v}^{T}\left(B_{H}\right)^{i-1}\mathbf{1}.$
5.3 Greedy Construction Algorithms
In order to get a clearer picture regarding the behavior of $n(H,g)$ we made several attempts to construct small girth $g$ graphs that cover $H_{23}$ by randomized greedy algorithms. Our starting point, Algorithm A, is an adaptation to $H_{23}$ of the Linial and Simkin [12] algorithm for regular graphs. The algorithms we describe are split into two categories:
  1. 1.
    Start with the initial graph $G_{0}$ being the cycle $C_{n}$ with vertex set $\{0,\ldots n-1\}$ where $i$ is connected to $i\pm 1\pmod{n}$. Keep adding edges, one at a time, until either the required graph is obtained, or there is no suitable edge to add. In step $k$, given the current graph $G_{k}$, let $P_{k}$ denote the graph of permissible new edges, where $V(P_{k})$ is the set of deficient $G_{k}$ vertices, $V(P_{k})=\{i:i\mbox{ is even and }\deg_{G_{k}}(i)=2\}$, and $E(P_{k})$ is the set of edges that can be added to $G_{k}$ without creating a cycle shorter than $g$. In step $k$ the algorithm randomly picks an edge $e$ in $E(P_{k})$ and adds it to $G_{k}$, where algorithms A,B and C differ only in the way they pick the edge $e$. For each $n,g$ pair, we perform multiple runs of the selected algorithm and check if at least one is successful.
  2. 2.
    Start with the initial graph $G_{0}$ being the 4 vertex clique $K_{4}$ minus an edge, depicted in Figure 3 (c). If at step $k$ we have $\mbox{girth}(G_{k})\geq g$ then stop. Otherwise randomly choose two edges $e,f\in E(G_{k})$, where the endpoints of each edge have different degrees, and obtain the graph $G_{k+1}$ by applying the operation shown in Figure 5. Both algorithms GD and GF fall into this framework and differ only in the way they pick $e,f$. For each value of $g$ we perform multiple runs of the selected algorithm and take the minimal size resulting graph.
    Figure 5: Transform a covering graph of $H$ by picking two edges $e=(u_{1},v_{1})$, $f=(u_{2},v_{2})$ with $\deg(u_{i})=3$ (marked by a square) and $\deg(v_{i})=2$ (marked by a circle). The new graph has 4 new vertices, 2 of degree two and 2 of degree three, and also covers $H$.
Algorithm A - Start with the cycle $C_{n}$ as above, where in step $k$ we pick $e\in E(P_{k})$ uniformly at random.
Algorithm B - The same, but in step $k$ we pick a uniformly random $u\in V(P_{k})$ and then pick $e=(u,v)\in E(P_{k})$ where $v$ is a uniformly random $P_{k}$ neighbor of $u$.
Algorithm C - The same, but in step $k$ we pick a uniformly random $u\in V(P_{k})$ from the vertices with smallest $\deg_{P_{k}}(u)$ and then pick $e=(u,v)\in E(P_{k})$ where $v$ is a uniformly random $P_{k}$ neighbor of $u$.
Algorithm GD - Start with $K_{4}$ minus an edge, where in each step one picks a uniformly random edge $e$ on a cycle of length smaller than $g$, and then choose a random $f$ from the edges that are most distant from $e$.
Algorithm GF - Start with $K_{4}$ minus an edge, where in each step one picks a uniformly random edge $e$ maximizing the the vector $c(e)=(c_{1}(e),c_{2}(e),\ldots,c_{g-1}(e))$ in lexicographic order, where $c_{l}(e)$ is the number of non-backtracking cycles of length $l$ in which $e$ participates. Given $e=(u_{1},v_{1})$ and $f=(u_{2},v_{2})$, define $D(e,f)=\mbox{dist}(v_{1},v_{2})+3$ and choose $f$ as follows: If $\max_{f}D(e,f)<g$ then choose a uniformly random $f$ maximizing $D(e,f)$. Otherwise, choose a uniformly random $f$ with $D(e,f)\geq g$ and maximal $c(f)$ in lexicographic order.
5.4 Result Summary
It should be noted that any graph of girth $g\geq 3$ is necessarily simple and therefore must have an even number of degree three vertices. Consequently, $n$ must be divisible by 4, so our results can always rounded up/down to the closest multiple of 4. Results are presented both as a graph in Figure 6 and as a table. For girth between 7 and 15, we improve upon the Moore bound by performing an exhaustive search. Also, for $g\in\{8,12,13,14,15\}$ an exhaustive search found smaller graphs than those found by the greedy algorithms.
Figure 6: Lower and upper bounds on $n(H_{23},g)$
# of Lower bounds Upper bounds
$g$ trials $n_{0}(H,g)$ best best A B C GD GF $n_{ES}(H,g)$
3 100k 4 4 4 4 4 4 4 4 -
4 100k 8 8 8 8 8 8 8 8 52
5 100k 8 8 8 8 8 8 8 8 84
6 100k 12 12 12 12 12 12 12 12 132
7 100k 16 20 20 20 20 20 20 20 200
8 100k 20 20 20 24 24 24 24 24 308
9 100k 24 28 28 28 28 28 28 28 476
10 100k 32 32 32 40 40 40 40 44 724
11 100k 40 48 48 48 48 48 48 48 1104
12 100k 48 52 52 76 72 60 72 72 1684
13 100k 60 76 76 104 104 80 100 92 2564
14 100k 76 88 88 144 144 104 140 128 3908
15 100k 96 104 144 248 232 168 208 176 5944
16 100k 116 116 232 348 324 240 284 232 9044
17 100k 144 144 332 528 492 352 404 332 13772
18 100k 176 176 452 816 732 508 568 452 20948
19 10k 224 224 640 1256 1172 776 796 640 31872
20 10k 272 272 916 1896 1780 1124 1144 916 48500
21 10k 340 340 1272 2928 2620 1668 1616 1272 73780
22 1k 412 412 1786 4432 3996 2484 2248 1786 112260
23 1k 520 520 2536 7524 6400 3872 3236 2536 170792
24 1k 628 628 3560 10964 9504 5848 4580 3560 259828
25 1k 792 792 5044 15880 14860 8568 6492 5044 395324
26 1k 960 960 7124 26048 20340 12524 9264 7124 601428
27 1k 1208 1208 10188 37656 29936 13088 10132 914992
28 1k 1456 1456 14488 58792 46900 18932 14488 1392084
29 1k 1836 1836 20704 98872 72940 27080 20680 2117860
30 1k 2220 2220 29784 129988 112016 38936 29712 3222084
We end with Figure 7, depicting the $(H,g)$-cages achieving the tight $n(H,g)$ results for $g\leq 10$.
Figure 7: $H$-cages for $g=3,4,\ldots,10$.
6 Conclusion & Open Problems
In this work we suggest alternatives, applicable to non-regular graphs, for the well studied problem of finding $n(k,g)$, which is the smallest size of a $k$-regular girth $g$ graph. Given a fixed base graph $H$, we define: (i) $n({\cal L}(H),g)=n(H,g)$ as the smallest size of a girth $g$ graph covering $H$, (ii) $n({\cal U}(H),g)$ as the smallest size of a girth $g$ graph with the same universal cover as $H$, and (iii) $n({\cal D}(H),g)$ as the smallest size of a girth $g$ graph with the same degree distribution as $H$. It should be noted that the second and third definitions offer a true generalization of $n(k,g)$, by taking any $k$-regular graph as $H$.
We prove that the first two definitions are equivalent up to a multiplicative constant, and we show that the well known Moore lower bound and Erdös-Sachs upper bound hold in the new setting as well, with the exponent base being $\rho(H)$ rather than k1k-1. For the 3rd definition we have a Moore bound result by Alon, Hoory and Linial [1]. However, there is no Erdös-Sachs result and the Moore bound, with exponent base $\Lambda(H)$, might be far from achievable as we conjecture is the case for $H_{23}$.
The question regarding the asymptotic behavior of $n(H,g)$, or $n(k,g)$ remains largely unanswered with a large gap between the lower and upper bounds. Constructions of better families of graphs, such as Lubotzky, Phillips and Sarnak [13], better graphs for small values of $g$ such as in this work for $n(H_{23},g)$ and the survey of $n(3,g)$ graphs by Exoo and Jajcay [6] point towards a conjecture that the truth is much closer to the Moore bound than what is currently know.
We end this work with a list of conjectures and open problems:
  1. 1.
    Conjecture: $n({\cal D}(H_{23}),g)\geq{\rho(H_{23})}^{(1+o(1))g/2}$
    Namely, the exponent base lower bound on the size of girth $g$ graphs where half the vertices have degree 2 and the other half have degree 3, can be improved from $\Lambda(H_{23})$ to $\rho(H_{23})$. An even bolder conjecture would be that graphs achieving the minimum cannot have adjacent degree 2 vertices.
  2. 2.
    Conjecture: The minimal possible $\rho$ for a graph in ${\cal D}(H_{23})$ is $\rho(H_{23})$.
    One should note that by Proposition 4.3 states that $\rho(G)>\Lambda(H_{23})$ for any graph $G\in{\cal D}(H_{23})$. However, it falls short of proving that $\rho(G)>\Lambda(H_{23})+\epsilon$ for some explicit $\epsilon>0$.
  3. 3.
    Problem: Prove that any of the greedy algorithms listed in Section 5 satisfies the asymptotic Erdös-Sachs upper bound of Theorem 1.1 whp.
    At least for Algorithm A, one may attempt to adapt the proof by Lineal and Simkin [12] to the non-regular case.
  4. 4.
    Problem: Prove that the Moore lower bound for $H_{23}$, given by (10) cannot hold as equality for large values of $g$, even when taking the divisibility constraint that $n(H_{23},g)$ must be divisible by four.
  5. 5.
    Problem: Provide a number theoretic construction of girth $g$ graphs covering some non-regular graph $H$ of size ${\rho(H)}^{(1+o(1))\cdot 3g/4}$.
Acknowledgements
I would like to thank my students Michael Shevelin, Tal Shamir, Almog Elbaz and Shaked Sharon for exploring greedy algorithms to construct small girth $g$ graphs covering $H_{23}$ and improving the exhaustive search for small values of $g$ as part of an undergrad project. Also, I would like to thank Dani Kotlar and Nati Linial for many fruitful discussions.